# Rule-based Data Selection for Large Language Models

### v2

@@ 3.2 Determinantal point process (DPP)
The optimal solution to this mathematical problem of selecting the most orthogonal subset of a set of vectors is NP-hard (Civril & Magdon-Ismail, 2007; Kulesza et al., 2012) but we use DPP sampling to provide a relatively good solution. The determinant point process (DPP) is a probabilistic model that describes the likelihood of selecting diverse subsets from a larger set (Macchi, 1975; Borodin & Olshanski, 2000). Mathematically, a DPP is defined by a kernel matrix that describes the similarities between elements in a set. The probability of selecting a particular subset is proportional to the determinant of the corresponding submatrix of this kernel matrix. Intuitively, subsets with highly similar items (leading to higher correlation in the submatrix) have smaller determinants and are thus less likely to be chosen.
@@
DPP Sampling. To sample a diverse subset using DPP, there are several existing algorithms (Hough et al., 2006; Kulesza et al., 2012; Tremblay et al., 2018) and the Python library DPPy (Gautier et al., 2019) implements some of them. The computation of the DPP sampling primarily hinges on the overhead of computing the inner product kernel matrix ${\bm{K}}$ and its eigendecomposition. In our case, ${\bm{K}}\in\mathbb{R}^{R\times R}$ and hence it requires $O(R^{3})$ time, where $R$ is the number of all rules. Nonetheless, we set $R=50$ in our experiments, therefore our DPP rule selection algorithm is extremely fast (typically within 0.1 seconds). Further details about DPP sampling algorithms and their time complexities can be found in Appendix A.2.
@@ 3.3 DPP rule-based rating algorithm
Figure 1: Pipeline for rule-based data rating and selection. Step 1. Use LLM to generate a comprehensive set of $R$ rules. Step 2. Rate a batch of $n$ data according to $R$ rules and form the score matrix ${\bm{S}}\in\mathbb{R}^{n\times R}$. Step 3. Select $r$ rules that correspond to the columns sampled by the DPP in the score matrix. Step 4. Rate the full dataset using the $r$ selected rules and (stochastically) select data with the highest averaged ratings. Step 5. Application of chosen data on downstream tasks, such as for LLM training.
The pipeline of our rule-based data selection method is illustrated in Figure 1 and comprises the following steps:
Step 1. Rule generation. We query GPT-4 to generate $R$ rules. In the prompt, we include the goal, the description of the source data, and the description of the downstream task to help GPT-4 generate relevant task-related rules.
Step 2. Rule-based rating: Recall the definitions in Section 3.1. We employ LLM, particularly Llama3-8B-Instruct (AI@Meta, 2024), to rate the batch data $\mathcal{B}$ according to $R$ rules, resulting in the matrix ${\bm{S}}\in\mathbb{R}^{n\times R}$.
@@
Note that the cost of generating $R$ rules is negligible, requiring just a single GPT-4 query, and the cost of obtaining the rating matrix ${\bm{S}}$ can be managed by adjusting the batch size $n$. The motivation to select a fixed small number of $r$ rules is driven by the computational costs associated with using LLMs for data rating and the need to maintain a consistent dimensionality for explaining data quality. These practical considerations lead us to treat $r$ as a hyperparameter. Discussions on the optimal choices of $r$ are explored in Section 4 and Appendix A.6.5.
@@
Step 4. Stochastic data selection: We extend the rating process to cover all data samples using the selected $r$ rules, expanding the rating matrix $\bar{{\bm{S}}}$ from $n\times r$ to $N\times r$. We then aggregate these fine-grained ratings by averaging across the $r$ columns of $\bar{{\bm{S}}}$, resulting in a score vector ${\bm{v}}=[v_{1},v_{2},\dots,v_{N}]$ that assigns a quality score to each of the $N$ samples.
Given the $N$ scores and a fixed budget of selecting $k$ samples for training, rather than choose the traditional top-$k$ approach, (selecting the $k$ highest scored samples), we adopt a stochastic sampling strategy, where we sample $k$ data points according to the distribution:
$p({\bm{x}}_{i})=\frac{e^{v_{i}}}{\sum_{j=1}^{N}e^{v_{i}}}$ (3)
for each data point ${\bm{x}}_{i}\in\mathcal{D}$. This stochastic data selection mechanism introduces greater diversity into the sampling process and is used in several other papers ((Wettig et al., 2024; Sachdeva et al., 2024)).
Step 5. Apply the selected data on given downstream tasks, such as for LLM pre-training and domain fine-tuning.
4 Evaluation A: Evaluating Against Ground Truth Ratings
We evaluate our method in two ways: A. by comparing the rating results against the ground truth rating of the dataset. Smaller deviations from the ground truth scores indicate better performance. Specifically, we rely on pairwise comparisons generated by GPT-4 and apply the Bradley-Terry model (Bradley & Terry, 1952) to compute $n$ scores, treating them as the ground truth. B. by training an LLM (Llama3-8B) with the selected data and assessing its performance through both general and domain-specific benchmarks. In this section, we present the first set of experiments (corresponding to Evaluation A), while the second set of experiments (based on Evaluation B) is discussed in Section 5. The low cost of Evaluation A enables us to explore various aspects such as the rule-size scaling law, different rating schemes (pairwise vs. single), and the impact of model sizes (Llama3-8B and Llama3-70B). These experiments provide preliminary evaluations of our method.
@@ 4.1 Experiments Setup
Datasets: We consider two datasets: CommonCrawl (Common Crawl, 2024), containing raw web-crawled data, and IMDB (Maas et al., 2011), a dataset of 50K movie reviews, representing general and domain-specific settings, respectively. For each dataset, we collect the first 50 examples and apply a pairwise comparison scheme for data rating (prompt templates are available in Appendix A.5.6), which requires comparison on 2,450 ordered pairs.
Ground truth scores: Ground truth scores are generated as follows: we prompt GPT-4 to compare each pair of data samples $(i,j)$ and then reversing the comparison for $(j,i)$. We only keep the pairs where both comparisons are consistent, filtering out cases where GPT-4 performs poorly. After filtering, approximately 1000 comparisons remain for CommonCrawl and 1800 for IMDB. From these outcomes, we calculate scores for the 50 samples using the Bradley-Terry model (Bradley & Terry, 1952) (details can be found in Appendix A.5.1).
Rating: Now with the ground truth scores, we use our rule-based approach to rate the same data. For each rule $i\in\{1,2,\dots,R\}$ ($R=50$), we employ Llama3-8B-Instruct as our comparison rater and similarly use the Bradley-Terry model to compute a score vector $S_{i}\in\mathbb{R}^{n}$ ($n$ is also 50 here), thereby forming the rating matrix ${\bm{S}}\in\mathbb{R}^{n\times R}$. Recall we denote $\bar{{\bm{S}}}$ as the submatrix of ${\bm{S}}$ containing $r$ columns indexed by the $r$ selected rules. To assess the rating results in $\bar{{\bm{S}}}$ against the ground truth, we compute the mean squared error (MSE):
@@
where $S_{GT}\in\mathbb{R}^{n}$ is the ground truth score vector and $\bar{S}_{j}$ is the $j$-th column of $\bar{{\bm{S}}}$. Furthermore, to establish comparative baselines, we implemented the same rating procedure (pairwise comparisons and score calculations via the Bradley-Terry model) using both the four designed rules in QuRating (see Wettig et al. (2024)) and a rule-free approach, referred to as the “NoRule” setting.
Our experiments in this section aim to address the following research questions: (Q1) Does greater rule diversity lead to more accurate ratings? (Q2) Does rule-based selection generally outperform rule-free methods? (Q3) How does our DPP-based rule selection compare to human-designed rules and ratings without rules? (Q4) Does DPP select better rules than randomly chosen ones? (Q5) How do different rating schemes and rater models impact the performance of our method?
@@ 4.2 Results
Correlation of $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$ (answer to Q1). For each $r\in\{1,2,\dots,50\}$, we sample $\min\{10000,\binom{50}{r}\}$ sets of indices of size $r$, which are used to choose rules and form $\bar{{\bm{S}}}$. We then calculate its rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$. We compute their Pearson correlation and observe positive values for both IMDB and CommonCrawl datasets (see Figures 2(a) and 2(b)). This confirms that higher rule diversity is positively correlated with the accuracy of rating results. In other words, the correlation or redundancy of rules is positively correlated with the error $\epsilon(\bar{{\bm{S}}})$.
Rule-based v.s. Rule-free (answer to Q2): We sample $10^{6}$ possible rule subsets with size $r$ from all 50 rules and calculate the corresponding MSE, comparing it to the MSE from the NoRule setting. The results in Figures 2(c) and 2(d) demonstrate that using rule-based rating is mostly guaranteed to give better results than rating without rules, no matter applied to general data like CommonCrawl or domain-specific data like IMDB. When compared to QuRating MSE, the results show that QuRating is outperformed by most randomly selected rule subsets, highlighting the limitations of human-designed rules.
@@
Figure 2: (a) and (b): Pearson correlation of the rule correlation $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$ for IMDB and CommonCrawl datasets respectively. (c) and (d): Distribution of MSE across $10^{6}$ random rule subsets with size $r$, for IMDB and CommonCrawl datasets respectively, with vertical lines representing the MSE values of QuRating and NoRule.
DPP v.s. QuRating v.s. NoRule (answer to Q3). For each $r\in\{1,2,\dots,50\}$, we use DPP to sample $r$ rules and conduct 100 trials. Then compare the averaged MSE against the MSEs from QuRating and NoRule, recording the winning rates of the DPP rules (see Figure 3(a)). For the IMDB dataset, we found that once $r$ reaches a certain threshold, DPP rules consistently achieve near-perfect winning rates against both NoRule and QuRating. Interestingly, for CommonCrawl, DPP underperforms QuRating when $r$ is too small or too large. This suggests that while QuRating rules are effective for general pre-training data, they lack the flexibility to adapt to other settings or domains.
DPP rules v.s. Randomly selected rules (answer to Q4). We compare DPP-selected rules with randomly selected rules of the same size $r$, evaluating both the rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$ for their corresponding score submatrices $\bar{{\bm{S}}}$. The results show that DPP consistently produces rules with lower correlation and MSE, regardless of the value $r$ (see Figures 3(b) and 3(c)). Another key observation is that the MSE for DPP rules increases when $r$ is either too small or too large, with the optimal $r$ falling somewhere in the middle. This matches our intuition: when $r$ is too small, there are too few rules to achieve sufficient rating diversity, and when $r$ is too large, rule redundancy can negatively affect the rating outcomes. In fact, this motivates our selection of $r=10$ in this paper.
@@
Figure 3: (a) Winning rate of DPP-selected rules compared to QuRating’s four rules and the NoRule setting, based on MSE across 100 DPP trials. (b) Comparison of rule correlation between DPP-selected and randomly selected rules, averaged across 100 trials. (c) Comparison of MSE between DPP-selected and randomly selected rules, averaged across 100 trials. Plots (a), (b), and (c) display results for the IMDB dataset, while (d), (e), and (f) for the CommonCrawl dataset.
Variations in rating schemes and rater models (answer to Q5). To verify that our method works across different rating schemes and rater models, we explored the following variations: 1. Pairwise v.s. individual rating. While the pairwise ratings provide more reliable comparisons, individual rating requires only $O(n)$ computation. We observed similar results as in Section 4.2 (see Appendix A.5.3). Notably, individual ratings on the IMDB dataset showed a Pearson correlation between rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$ of up to 0.6, and the winning rates show that DPP significantly outperforms both QuRating and the NoRule. 2. Llama3-8B v.s. Llama3-70B. We tested the influence of rater model capability by switching to Llama3-70B (instruction-tuned version), using the individual rating scheme on IMDB. The results are similar to earlier and we also noted a high Pearson correlation (over 0.6) between rule correlation and MSE, along with a high winning rate of DPP compared to QuRating and NoRule. Furthermore, randomly selected rules perform significantly better than both QuRating and NoRule. See Appendix A.5.4 for further details.
5 Evaluation B: Data Selection for LLM Fine-tuning
In this section, we follow the pipeline outlined in Section 3.3 and conduct experiments based on Evaluation B, where we train an LLM (Llama3-8B) using the selected data and assess its performance. This setup closely reflects real-world applications of LLM data selection. We benchmark our method against several baselines, such as uniform sampling, direct rating without rules, QuRating rules (Wettig et al., 2024), and DSIR (Xie et al., 2024) (a commonly used baseline for LLM data selection). We condcut experiments in this section to explore the following research questions: (Q1) How does data selected by rule-based methods enhance model fine-tuning compared to rule-free methods? (Q2) How do the rules generated by our automated framework compare to human-designed rules? (Q3) How does DPP rule selection perform compared to random rule selection?
@@ 5.1 Experiments Setup
Evaluation Benchmarks. To systematically evaluate the effectiveness of our framework, we use following benchmarks: For experiments on general continued pre-training, we utilize ARC-Easy (Yadav et al., 2019), ARC-Challenge (Yadav et al., 2019), Winogrande (Sakaguchi et al., 2021), MMLU(Hendrycks et al., 2020), and SST-2 (Socher et al., 2013). Then we employ domain-specific datasets to do fine-tuning: For IMDB, we use the IMDB sentiment analysis dataset (Maas et al., 2011). For Code, we use benchmarks for code generation, including HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), Multiple-py and Multiple-cpp (Cassano et al., 2022). For Math and Medical domains, we choose subsets from MMLU corresponding to Math subject and Medical subject respectively. More details about these benchmarks are summarized in Appendix A.6.4.
Data Source. SlimPajama is a large, deduplicated, multi-corpus open-source dataset specifically designed for training large language models (Cerebras Systems, 2023). We randomly sampled 1 million data points (around 1 billion tokens) from SlimPajama as our initial data source $\mathcal{D}$. From this pool, we employ our selection methods to choose data for training.

### v3

@@ 3.2 Determinantal point process (DPP)
The optimal solution to this mathematical problem of selecting the most orthogonal subset of a set of vectors is NP-hard (civril2007finding; kulesza2012determinantal) but we use DPP sampling to provide a relatively good solution. The determinant point process (DPP) is a probabilistic model that describes the likelihood of selecting diverse subsets from a larger set (macchi1975coincidence; borodin2000distributions). Mathematically, a DPP is defined by a kernel matrix that describes the similarities between elements in a set. The probability of selecting a particular subset is proportional to the determinant of the corresponding submatrix of this kernel matrix. Intuitively, subsets with highly similar items (leading to higher correlation in the submatrix) have smaller determinants and are thus less likely to be chosen.
@@
DPP Sampling. To sample a diverse subset using DPP, there are several existing algorithms (hough2006determinantal; kulesza2012determinantal; tremblay2018optimized) and the Python library DPPy (GPBV19) implements some of them. The computation of the DPP sampling primarily hinges on the overhead of computing the inner product kernel matrix $\bm{K}$} and its eigendecomposition. In our case, ${\bm{K}}\in\mathbb{R}^{R\times R}$ and hence it requires $O(R^{3})$ time, where $R$ is the number of all rules. Nonetheless, we set $R=50$ in our experiments, therefore our DPP rule selection algorithm is extremely fast (typically within 0.1 seconds). Further details about DPP sampling algorithms and their time complexities can be found in Appendix E.
@@ 3.3 DPP rule-based rating algorithm
Figure 1: Pipeline for rule-based data rating and selection in five steps (detailed in Section 3.3)
Our rule-based data selection framework proceeds in five steps (Figure 1):
Step 1. Rule generation. We prompt GPT-4 with a task description and source data context to generate $R$ candidate rules. These rules are then manually refined for clarity and relevance.
Step 2. Rule-based rating: We employ Llama3-8B-Instruct (llama3modelcard), to rate the batch data $\mathcal{B}$ according to $R$ rules, resulting in the score matrix ${\bm{S}}\in\mathbb{R}^{n\times R}$.
@@
Note that the cost of generating $R$ rules is negligible, requiring just a single GPT-4 query, and the cost of obtaining the rating matrix $\bm{S}$} can be managed by adjusting the batch size $n$. The motivation to select a fixed small number of $r$ rules is driven by the computational costs associated with using LLMs for data rating and the need to maintain a consistent dimensionality for explaining data quality. These practical considerations lead us to treat $r$ as a hyperparameter. Discussions on the optimal choices of $r$ are explored in Section 4 and Appendix I.5.
@@
Step 4. Stochastic data selection: We extend the rating process to cover all data samples using the selected $r$ rules, expanding the rating matrix $\bar{{\bm{S}}}$ from $n\times r$ to $N\times r$. We then aggregate these fine-grained ratings by averaging across the $r$ columns of $\bar{{\bm{S}}}$, resulting in a score vector ${\bm{v}}=[v_{1},v_{2},\dots,v_{N}]^{\top}$ that assigns a quality score to each of the $N$ samples.
Given the $N$ scores and a fixed budget of selecting $k$ samples for training ($k=20000$ in Section 5), rather than choose the traditional top-$k$ approach, (selecting the highest scored samples), we adopt a stochastic sampling strategy, where we sample $k$ data points according to the distribution:
$p({\bm{x}}_{i})=\frac{e^{v_{i}/\tau}}{\sum_{j=1}^{N}e^{v_{i}/\tau}}$ (4)
for each data point ${\bm{x}}_{i}\in\mathcal{D}$, and $\tau$ is the “temperature” balancing between top-$k$ and uniform sampling (we use $\tau=1$ by default). This stochastic data selection mechanism introduces greater diversity into the sampling process and is used in several other works (wettig2024qurating; sachdeva2024train).
Step 5. Apply the selected data on given downstream tasks, such as for LLM domain fine-tuning.
4 Preliminary Experiments: Evaluating Against Ground Truth Ratings
We evaluate our method through two complementary approaches: (A) comparing rule-based ratings against ground truth human ratings, where smaller deviations from human scores indicate better performance, and (B) training LLMs (Pythia-1B and Llama3-8B) on the selected data and evaluating their downstream performance on domain-specific benchmarks. This section focuses on the first evaluation setup (A), while the full-scale LLM training experiments (B) are presented in Section 5. Because Evaluation A avoids expensive LLM training, it enables a broader exploration of factors such as the effect of rule set size and the influence of model scale (Llama3-8B vs. Llama3-70B). These experiments offer deeper insights into the robustness and effectiveness of our approach.
@@ 4.1 Experiments Setup
Datasets and ground-truth ratings: We use four datasets spanning different domains: StanfordNLP-IMDB (maas-EtAl:2011:ACL-HLT2011) (IMDB), MedQA (jin2021disease) (Medical), GSM8K (cobbe2021training) (Math), and MBPP (austin2021program) (Code). For each dataset, we randomly sample 300 examples and ask five human annotators to assign quality scores, with the average scores treated as ground truth (GT). In this section, we present results on IMDB as a representative example; results on the remaining domains are provided in Appendix H.2.
Rule-based rating: We apply our rule-based approach to rate the same data. For each of the $R=50$ rules, we obtain ratings from two LLMs: Llama3-8B-Instruct and Llama3-70B-Instruct, to evaluate the effect of rater capability. Each rule $i$ produces a score vector $S_{i}\in\mathbb{R}^{n}$ over $n=300$ samples, forming a full rating matrix ${\bm{S}}\in\mathbb{R}^{n\times R}$. We then select a subset of $r$ rules, forming a submatrix $\bar{{\bm{S}}}\in\mathbb{R}^{n\times r}$. To evaluate how well this rule subset aligns with ground-truth scores $S_{\text{GT}}\in\mathbb{R}^{n}$, we compute the mean squared error (MSE):
@@
where $\bar{S}_{j}$ is the $j$-th column of $\bar{{\bm{S}}}$. As baselines, we include: (1) the four human-designed rules in QuRating (wettig2024qurating), and (2) a rule-free method where Llama3 directly produces an overall quality score (“NoRule”).
Our experiments in this section aim to explore the following research questions:
  • (Q1) Does greater rule diversity lead to more accurate ratings?
  • (Q2) Does rule-based selection generally outperform rule-free methods?
  • (Q3) Is the rating quality based on our DPP-selected rules better than that based on human-created rules and NoRule setting?
  • (Q4) Does DPP select better rules than randomly chosen ones?
  • (Q5) Can the method possibly generalize well to pre-train datasets?
@@ 4.2 Results
Correlation of $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$ (answer to Q1). For each rule subset size $r\in\{1,2,\dots,50\}$, we sample $\min\{10000,\binom{50}{r}\}$ sets of indices of size $r$, which are used to choose rules and form $\bar{{\bm{S}}}$. We then calculate its rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$, as well as the Pearson correlation between them. The Pearson correlation is surprisingly strong across both rating model (Llama3 8B and 70B) (Figures 2(a) and 2(b)). This confirms that higher rule diversity is positively correlated with the accuracy of rating results. In other words, the highly correlated or redundant rules can potentially lead to non-robust ratings.
Rule-based v.s. Rule-free (answer to Q2): We sample $10^{6}$ random rule subsets of size $r$ and compare their MSEs against that of the NoRule baseline. Figures 2(c) and 2(d) show that rule-based methods consistently achieve lower MSE. Interestingly, even randomly chosen domain-specific rules outperform the fixed QuRating rules, illustrating the limitations of static, general-purpose rule sets.
@@
Figure 2: (a) and (b): Pearson correlation of the rule correlation $\rho(\bar{{\bm{S}}})$ and the MSE $\epsilon(\bar{{\bm{S}}})$, using Llama3 8B and 70B raters respectively. (c) and (d): Distribution of MSE from $10^{6}$ possible rule subsets with size $r$, using Llama3 8B and 70B raters respectively, where two vertical lines represent the MSE values of QuRating and NoRule.
DPP v.s. QuRating v.s. NoRule (answer to Q3). Using DPP, we select $r$ rules and repeat this for 100 trials. We report both the average MSE and the winning rate compared to QuRating and NoRule (Figures 3(a), 3(d)). Results show that once $r$ exceeds a modest threshold, DPP-selected rules almost always outperform both baselines.
@@
Figure 3: (a) Winning rate of DPP-selected rules compared to QuRating’s four rules and the NoRule setting respectively, based on MSE across 100 DPP trials. (b) Comparison of rule correlation $\rho$ between DPP-selected and randomly selected rules, averaged across 100 trials. (c) Comparison of MSE between DPP-selected and randomly selected rules, averaged across 100 trials. Plots (a), (b), and (c) display results using Llama3-8B rater, while (d), (e), and (f) for the Llama3-70B rater.
DPP rules v.s. Randomly selected rules (answer to Q4). We compare DPP-selected rule subsets to randomly sampled ones of the same size, evaluating both the rule correlation $\rho(\bar{{\bm{S}}})$ and MSE $\epsilon(\bar{{\bm{S}}})$ for their corresponding score submatrices $\bar{{\bm{S}}}$. As shown in Figures 3(b) and 3(c), DPP consistently selects rules with lower correlation and MSE, regardless of the value $r$. Notably, MSE increases when $r$ is too small or too large, which matches our intuition: when $r$ is too small, there are too few rules to cover all important data quality aspects for rating, and when $r$ is too large, rule redundancy can negatively affect the rating outcomes. Therefore, we use $r=10$ in subsequent experiments.
Apply to pre-training datasets (answer to Q5). So far, our experiments have focused on domain-specific datasets. We also applied the rule-based method to the CommonCrawl dataset (commoncrawl), a general-purpose web corpus commonly used for LLM pre-training. We find that the plots of rule correlation show similar patterns, and the Pearson correlation between $\rho$ and $\epsilon$ is consistently positive. However, the trends and Pearson values are less pronounced than in domain-specific settings. This may be because domain-specific datasets support more concrete and well-defined quality rules, whereas quality assessment in general pretraining data tends to be more ambiguous and less structured. Detailed results and analysis are provided in Appendix H.3.
5 Experiments: Data Selection for LLM Fine-tuning
We now evaluate our full framework in a realistic setting by following the pipeline in Section 3.3 to fine-tune LLMs (Pythia-1B and Llama3-8B) on selected data, then measure their downstream performance. This setup reflects practical applications of LLM data selection in domain-specific fine-tuning.
@@ 5.1 Experiments Setup
Evaluation Benchmarks. We assess performance across four domains using standard benchmarks: IMDB: Sentiment classification using the IMDB dataset (maas-EtAl:2011:ACL-HLT2011), Code: Code generation evaluated on HumanEval (chen2021evaluating), MBPP (austin2021program), and Multiple-py/cpp (cassano2022multipl), Math/Medical: Relevant subsets of MMLU (hendrycks2020measuring). We set the inference temperature to 0, unless the the benchmark protocol itself requires top-$k$ sampling. Full benchmark details are provided in Appendix I.4.
Data Source. We use SlimPajama (CerebrasSlimPajama2023), a large, deduplicated multi-source corpus for LLM training. From it, we randomly sample 1M examples (around 1B tokens) as our raw dataset $\mathcal{D}$, from which we select a high-quality subset of k=20,000k=20{,}000 samples for model training.